# Federated Hierarchical Reinforcement Learning for Adaptive Traffic Signal Control

### v1

@@ Abstract
Multi-agent reinforcement learning (MARL) has shown promise for adaptive traffic signal control (ATSC), enabling multiple intersections to coordinate signal timings in real time. However, in large-scale settings, MARL faces constraints due to extensive data sharing and communication requirements. Federated learning (FL) mitigates these challenges by training shared models without directly exchanging raw data, yet traditional FL methods such as FedAvg struggle with highly heterogeneous intersections. Different intersections exhibit varying traffic patterns, demands, and road structures, so performing FedAvg across all agents is inefficient. To address this gap, we propose Hierarchical Federated Reinforcement Learning (HFRL) for ATSC. HFRL employs clustering-based or optimization-based techniques to dynamically group intersections and perform FedAvg independently within groups of intersections with similar characteristics, enabling more effective coordination and scalability than standard FedAvg. Our experiments on synthetic and real-world traffic networks demonstrate that HFRL not only outperforms both decentralized and standard federated RL approaches but also identifies suitable grouping patterns based on network structure or traffic demand, resulting in a more robust framework for distributed, heterogeneous systems.
@@ 1 INTRODUCTION
Traffic congestion has become a pressing issue in urban areas, exacerbated by the continuous growth in the number of vehicles. In 2023, drivers in New York City lost an average of $101$ hours due to traffic congestion, causing approximately \$9.1 billion in lost productivity (New York Post, 2024). Adaptive Traffic Signal Control (ATSC) systems have emerged as a promising solution to improve the efficiency of traffic flow by coordinating signal operations across multiple intersections in real time. By dynamically adjusting the traffic signal phase, ATSC aims to minimize delays and congestion, providing a more responsive traffic management system. Optimization methods are widely used in ATSC (Mohebifard et al., 2019; Mohebifard and Hajbabaie, 2019; Yin et al., 2023; Zheng et al., 2024).
@@
Recent years have seen a growing amount of literature that applies learning methods for ATSC, with the aim of learning the stochastic and unknown dynamics of traffic environments. Reinforcement learning (RL), in particular, has been widely applied to ATSC (Li et al., 2024; Mo et al., 2022; Zhang et al., 2024; Aslani et al., 2017; Wang et al., 2024). Traffic lights are modeled as intelligent agents that cooperate or optimize predefined traffic metrics in a decentralized manner while learning traffic dynamics. Specifically, RL-based approaches can dynamically adapt to real-time traffic conditions, making them well suited to handle variations in flow rates and unexpected congestion patterns. Moreover, these methods often outperform traditional rule-based solutions by continuously learning and refining control strategies based on ongoing feedback from the traffic environment (Wei et al., 2019; Wang et al., 2024).
@@
However, applying RL to ATSC presents several challenges. One potential concern arises when RL models rely on fine-grained data such as vehicle positions, velocities, or trajectories, because collecting and storing such information can raise privacy issues (Wang et al., 2018). While many RL-ATSC approaches rely solely on aggregated traffic counts or intersection-level flow data - reducing privacy risks - some recent methods incorporate detailed trajectory data to capture finer spatiotemporal dynamics (Essa and Sayed, 2020; Yazdani et al., 2023). Although these finer-grained inputs can improve RL-ATSC results, their collection and potential sharing among intersections still require careful consideration of how to protect the information of individual drivers. Implementing data anonymization and secure data sharing protocols can help mitigate these privacy concerns while still enabling advanced RL-ATSC methods to leverage detailed traffic data. Another potential issue in multi-agent reinforcement learning (MARL) is that sharing information between intersections can create a substantial communication burden (Wang et al., 2020, 2021). Consequently, there are an increasing number of studies that have employed Federated Reinforcement Learning (FRL) to ATSC (Wang et al., 2020; Fu and Di, 2023; Bao et al., 2023). FRL ensures vehicle privacy by keeping raw data local at each intersection and allowing each intersection to only share aggregated model parameters. It also reduces communication overhead by transmitting far less data compared to systems that require the exchange of detailed traffic information. In FRL for ATSC, the FedAvg (McMahan et al., 2017) algorithm is predominantly used (Ye et al., 2021; Fu and Di, 2023), allowing agents to send their model weights to the central server periodically and receive the averaged weights from other agents.
@@
FRL has emerged as a promising approach to address privacy and communication challenges in MARL for ATSC. However, intersections in large cities often exhibit significant heterogeneity in traffic volumes, demand patterns, and road geometries. Furthermore, the mix of pedestrians, micromobility, and motor vehicles can vary significantly between different neighborhoods, reflecting various demands of traffic (Zhou and Zhu, 2007). On the supply side, traffic management strategies, such as highway ramp metering and tiered traffic signal systems, illustrate a hierarchical structure, with major and minor roads serving distinct functions in regulating the overall traffic flow (Han et al., 2020). A single, uniformly applied approach like FedAvg may struggle to account for these variations, potentially resulting in suboptimal performance in large-scale networks.
@@ 1.1 Contribution
  • High Accuracy Despite Decentralized Training: By introducing FedFomoLight and FedClusterLight, we personalize local agents within the HFRL framework to accommodate varying traffic conditions. Remarkably, even in a decentralized setting, our method outperforms classical centralized approaches (e.g., FedAvg) in terms of travel time and waiting time.
@@ 2.1 MARL for Traffic Signal Control
RL is a branch of machine learning. In ATSC, RL-based methods do not depend on traffic models, but learn from past experience and adjust traffic signal phases according to global or local environments (Mo et al., 2022; Wang et al., 2024). MARL methods (Xu et al., 2021; Soleimany et al., 2023; Yu et al., 2023) surpass traditional single-agent approaches in urban ATSC by deploying individual agents at intersections to make decisions independently using local data. Typically, there are two types of MARL schemes: centralized and decentralized (Song et al., 2024; Wei et al., 2019; Wang et al., 2024). Centralized MARL uses a central controller to oversee the entire traffic network, making decisions for all traffic signals simultaneously. In contrast, decentralized MARL allows each traffic signal to operate as an independent agent, reducing communication overhead and adapting more effectively to dynamic conditions, making it ideal for large-scale urban areas (Li et al., 2024; Mo et al., 2022; Zhang et al., 2024; Aslani et al., 2017; Wang et al., 2024).
@@ 2.2 Federated Learning and Traffic Signal Control
FL is a distributed collaborative learning framework that enables multiple agents to train a shared machine learning model without exchanging raw data, but by sharing models’ weights. Initially introduced by Google (McMahan et al., 2017) to predict user text input on Android devices, FL uses local datasets on individual devices while maintaining data privacy and security. This approach is particularly advantageous in scenarios where privacy concerns and large-scale distributed systems are present, as it mitigates the need for central data pooling.
@@
Wang et al. (2020) Introduces two MARL control modes and a FRL framework to solve the ATSC problem. It focuses on improving robustness and computational efficiency by distributing control across multiple agents, allowing local optimization at each intersection. A2C MARL with independent and joint control mode, as well as an FL framework
Ye et al. (2021) Presents a cloud-based FRL framework that coordinates decentralized agent training for optimizing multi-intersection traffic control. Uses FL-like gradient aggregation and parameter sharing to enhance training convergence and enables knowledge sharing among agents. A2C MARL with FL
Hudson et al. (2022) Introduces a SEAL framework to represent traffic conditions across diverse intersections and proposes a FRL approach for training traffic light decision-making policies. Improves the reward-communication trade-off, reducing communication costs by up to 36.24% while maintaining only a 2.11% performance decrease compared to centralized training. PPO MARL with FL
Fu and Di (2023) Proposes an open-source simulation framework that integrates SUMO-Gym with Ray’s Rllib for FRL to solve the ATSC problem. Introduces the first A3C-based FL framework for ATSC and tests the approach on real-world road networks, calibrating traffic volumes to analyze communication costs. A3C MARL with FL
Soleimany et al. (2023) Introduces a hierarchical learning model to address scalability in traffic signal management. Uses profilization-based learning with FRL to process data locally at node and intersection levels, optimizing traffic flow while maintaining source-to-destination delays (SDDs) with deep RL. DRL RL on intersection level with FL on cloud level
Bao et al. (2023) Proposes an FRL method for TSC. Introduces a unified state representation and action selection method to handle diverse intersection structures, using partial aggregation and fine-tuning to preserve local specificity. DQN Deep Reinforcement Learning with FL
@@ 2.3 Heterogeneous Federated Learning
Existing FL approaches predominantly assume that all clients share a uniform network structure or similar data distributions. However, practical scenarios often exhibit substantial differences among clients in terms of data distributions, model structures, communication networks, and edge devices (Ye et al., 2023). Hierarchy naturally arises due to asymmetry in demand and supply. For instance, intersections near schools may experience different peak traffic hours compared to those near sports arenas. To address FL heterogeneity, several state-of-the-art methods have been proposed. These methods, known as HFL, are categorized into data-level, model-level, and server-level approaches (Ye et al., 2023). Among these approaches, it is crucial to identify one that can account for the heterogeneity of different intersections by dynamically grouping similar ones and training models that reflect their specific characteristics.
@@ 2.3.1 Data-level Methods
Data-level methods focus primarily on tasks related to data management, including data augmentation and anonymization techniques. Significant contributions in this area include (Li et al., 2020) (Li and Wang, 2019) (Yoon et al., 2021). While these efforts are aimed at mitigating statistical heterogeneity between clients and enhancing data privacy, they may not be ideally suited for grouping and training models that accurately reflect the unique characteristics of diverse datasets.
@@ 2.3.2 Model-level Methods
Model-level methods focus on operations designed specifically for the model layer, including sharing partial structures and model optimization. Recent research on $\Psi$-Net (Yu et al., 2022) and QAvg &\&& PAvg (Jin et al., 2022) optimizes policy performance in diverse environments without data sharing, although mainly for tabular methods. Per-FedAvg (Fallah et al., 2020) and FedPer (Arivazhagan et al., 2019) focus on personalized federated learning, primarily for supervised learning environments. Related work in FedFormer (Hebert et al., 2022) and the KT-pFL framework (Zhang et al., 2021) enhances personalization through innovative model structures but with potential limitations in direct policy updates.
@@ 2.3.3 Server-level Methods
Server-level methods focus on the central server’s engagement, including tasks such as client selection, grouping, and coordination. Related works include deep reinforcement learning (DRL) on the heterogeneous training intensity assignment problem for FL (Zeng et al., 2023) and RL client grouping (Pang et al., 2021), where both focus on dynamic adjustment via RL. Other works include FedCor (Tang et al., 2022) and QAvg &\&& PAvg (Jin et al., 2022).
Cluster-based and optimization-based approaches further enhance client coordination and model personalization at the server-level. Cluster-based methods infer the heterogeneity of the data distribution using model information, assigning clients to clusters for model aggregation. Examples include research on FedCHAR &\&& FedCHAR-DC (Li et al., 2023), the Federated Multitask Learning (FMTL) framework (Sattler et al., 2020), and the Iterative Federated Clustering Algorithm (IFCA) (Ghosh et al., 2020). Other clustering methods, such as hierarchical clustering (Briggs et al., 2020), EM clustering (Long et al., 2023), and KMeans (Ghosh et al., 2019), help FL converge faster with lower communication costs. In optimization-based approaches, client aggregation is often formulated as an integer programming problem to minimize objective functions. Such methods include Ditto (Li et al., 2021), the APPLE framework (Luo and Wu, 2022), and FedFomo (Zhang et al., 2020).
@@
$\mathcal{L}$ set of local loss functions
@@ 3.1 MARL
Formally, the ATSC on a road network problem is modeled as a Markov game defined by the tuple $\bigl{(}N,S,O,A,P,R,\gamma\bigr{)}$, where $N$ is the number of agents (intersections), and $S,O,A,P,R,\gamma$ denote the state space, joint observation space, action space, state transition functions, reward functions, and discount factor, respectively. Since each agent has only partially observability of the global state, the Markov game is formulated as a partially observable Markov decision process (POMDP). At each time step, each agent chooses an action based on its local policy, transitioning the environment to a new state, and generating a local reward. This cycle repeats until a terminal condition is reached, and agents refine their policies to maximize their individual rewards, thereby improving overall network performance.
@@
    State ($S$). For a traffic light $n$ at time $\tau$, the local state includes lane occupancy $\bigl{(}o_{n}^{\tau}\bigr{)}$, queue length $\bigl{(}q_{n}^{\tau}\bigr{)}$, average speed $\bigl{(}\overline{u}_{n}^{\tau}\bigr{)}$, and the phase ratios $\bigl{(}ph_{n}^{\tau}\bigr{)}$ for the relevant signal phases (e.g., green, yellow, red).
@@
    Action ($A$). To ensure consistency across agents in the federated learning setting, we define a binary action space $a_{n}^{\tau}\in\{0,1\}$. Here, $a_{n}^{\tau}=1$ indicates that traffic light $n$ will attempt to switch to the next phase at time $\tau$, while $a_{n}^{\tau}=0$ means that it remains in its current phase.

### v2

@@ Abstract
Multi-agent reinforcement learning (MARL) has shown promise for adaptive traffic signal control (ATSC), enabling multiple intersections to coordinate signal timings in real time. However, in large-scale settings, MARL faces constraints due to extensive data sharing and communication requirements. Federated learning (FL) mitigates these challenges by training shared models without directly exchanging raw data, yet traditional FL methods such as FedAvg struggle with highly heterogeneous intersections. Different intersections exhibit varying traffic patterns, demands, and road structures, so performing FedAvg across all agents is inefficient. To address this gap, we propose Hierarchical Federated Reinforcement Learning (HFRL) for ATSC. HFRL employs clustering-based or optimization-based techniques to dynamically group intersections and perform FedAvg independently within groups of intersections with similar characteristics, enabling more effective coordination and scalability than standard FedAvg.Our experiments on synthetic and real-world traffic networks demonstrate that HFRL consistently outperforms decentralized and standard federated RL approaches, and achieves competitive or superior performance compared to centralized RL as network scale and heterogeneity increase, particularly in real-world settings. The method also identifies suitable grouping patterns based on network structure or traffic demand, resulting in a more robust framework for distributed, heterogeneous systems.
@@ 1 INTRODUCTION
Traffic congestion has become a pressing issue in urban areas, exacerbated by the continuous growth in the number of vehicles. In 2023, drivers in New York City lost an average of $101$ hours due to traffic congestion, causing approximately \$9.1 billion in lost productivity (NYPostTraffic2024). Adaptive Traffic Signal Control (ATSC) systems have emerged as a promising solution to improve the efficiency of traffic flow by coordinating signal operations across multiple intersections in real time. By dynamically adjusting the traffic signal phase, ATSC aims to minimize delays and congestion, providing a more responsive traffic management system. Optimization methods are widely used in ATSC (mohebifard2019cooperative; mohebifard2019optimal; yin2023real; zheng2024integrated).
@@
Recent years have seen a growing amount of literature that applies learning methods for ATSC, with the aim of learning the stochastic and unknown dynamics of traffic environments. Reinforcement learning (RL), in particular, has been widely applied to ATSC (li2024cooperative; MO2022103728; zhang2024survey; aslani2017adaptive; wang2024large). Traffic lights are modeled as intelligent agents that cooperate or optimize predefined traffic metrics in a decentralized manner while learning traffic dynamics. Specifically, RL-based approaches can dynamically adapt to real-time traffic conditions, making them well suited to handle variations in flow rates and unexpected congestion patterns. Moreover, these methods often outperform traditional rule-based solutions by continuously learning and refining control strategies based on ongoing feedback from the traffic environment (wei2019presslight; wang2024large).
@@
However, applying RL to ATSC presents several challenges. One potential concern arises when RL models rely on fine-grained data such as vehicle positions, velocities, or trajectories, because collecting and storing such information can raise privacy issues (wang2018privacy). While many RL-ATSC approaches rely solely on aggregated traffic counts or intersection-level flow data - reducing privacy risks - some recent methods incorporate detailed trajectory data to capture finer spatiotemporal dynamics (essa2020self; yazdani2023intelligent). Although these finer-grained inputs can improve RL-ATSC results, their collection and potential sharing among intersections still require careful consideration of how to protect the information of individual drivers. Implementing data anonymization and secure data sharing protocols can help mitigate these privacy concerns while still enabling advanced RL-ATSC methods to leverage detailed traffic data. Another potential issue in multi-agent reinforcement learning (MARL) is that sharing information between intersections can create a substantial communication burden (wang2020large; wang2021adaptive). Consequently, there are an increasing number of studies that have employed Federated Reinforcement Learning (FRL) to ATSC (wang2020adaptive; fu2023federated; bao2023scalable). FRL ensures vehicle privacy by keeping raw data local at each intersection and allowing each intersection to only share aggregated model parameters. It also reduces communication overhead by transmitting far less data compared to systems that require the exchange of detailed traffic information. In FRL for ATSC, the FedAvg (mcmahan2017communication) algorithm is predominantly used (ye2021fedlight; fu2023federated), allowing agents to send their model weights to the central server periodically and receive the averaged weights from other agents.
@@
FRL has emerged as a promising approach to address privacy and communication challenges in MARL for ATSC. However, intersections in large cities often exhibit significant heterogeneity in traffic volumes, demand patterns, and road geometries. Furthermore, the mix of pedestrians, micromobility, and motor vehicles can vary significantly between different neighborhoods, reflecting various demands of traffic (jing2007hierarchy). On the supply side, traffic management strategies, such as highway ramp metering and tiered traffic signal systems, illustrate a hierarchical structure, with major and minor roads serving distinct functions in regulating the overall traffic flow (han2020hierarchical). A single, uniformly applied approach like FedAvg may struggle to account for these variations, potentially resulting in suboptimal performance in large-scale networks.
@@ 1.1 Contribution
  • High Accuracy Despite Decentralized Training: By introducing FedFomoLight and FedClusterLight, we personalize local agents within the HFRL framework to accommodate varying traffic conditions. Our methods achieve performance comparable to centralized approaches in small synthetic networks, and surpass centralized and classical FRL methods as network heterogeneity and scale increase, especially in real-world traffic networks.
@@ 2.1 MARL for Traffic Signal Control
RL is a branch of machine learning. In ATSC, RL-based methods do not depend on traffic models, but learn from past experience and adjust traffic signal phases according to global or local environments (MO2022103728; wang2024large). MARL methods (xu2021hierarchically; soleimany2023hierarchical; yu2023decentralized) surpass traditional single-agent approaches in urban ATSC by deploying individual agents at intersections to make decisions independently using local data. Typically, there are two types of MARL schemes: centralized and decentralized (song2024cooperative; wei2019presslight; wang2024large). Centralized MARL uses a central controller to oversee the entire traffic network, making decisions for all traffic signals simultaneously. In contrast, decentralized MARL allows each traffic signal to operate as an independent agent, reducing communication overhead and adapting more effectively to dynamic conditions, making it ideal for large-scale urban areas (li2024cooperative; MO2022103728; zhang2024survey; aslani2017adaptive; wang2024large).
@@ 2.2 Federated Learning and Traffic Signal Control
FL is a distributed collaborative learning framework that enables multiple agents to train a shared machine learning model without exchanging raw data, but by sharing models’ weights. Initially introduced by Google (mcmahan2017communication) to predict user text input on Android devices, FL uses local datasets on individual devices while maintaining data privacy and security. This approach is particularly advantageous in scenarios where privacy concerns and large-scale distributed systems are present, as it mitigates the need for central data pooling.
@@
wang2020adaptive Introduces two MARL control modes and a FRL framework to solve the ATSC problem. It focuses on improving robustness and computational efficiency by distributing control across multiple agents, allowing local optimization at each intersection. A2C MARL with independent and joint control mode and FL
ye2021fedlight Presents a cloud-based FRL framework that coordinates decentralized agent training for optimizing multi-intersection traffic control. Uses FL-like gradient aggregation and parameter sharing to enhance training convergence and enables knowledge sharing among agents. A2C MARL with FL
hudson2022smart Introduces a SEAL framework to represent traffic conditions across diverse intersections and proposes a FRL approach for training traffic light decision-making policies. Improves the reward-communication trade-off, reducing communication costs by up to 36.24% while maintaining only a 2.11% performance decrease compared to centralized training. PPO MARL with FL
fu2023federated Proposes an open-source simulation framework that integrates SUMO-Gym with Ray’s Rllib for FRL to solve the ATSC problem. Introduces the first A3C-based FL framework for ATSC and tests the approach on real-world road networks, calibrating traffic volumes to analyze communication costs. A3C MARL with FL
soleimany2023hierarchical Introduces a hierarchical learning model to address scalability in traffic signal management. Uses profilization-based learning with FRL to process data locally at node and intersection levels, optimizing traffic flow while maintaining source-to-destination delays (SDDs) with deep RL. DRL RL on intersection level with FL on cloud level
bao2023scalable Proposes an FRL method for TSC. Introduces a unified state representation and action selection method to handle diverse intersection structures, using partial aggregation and fine-tuning to preserve local specificity. DQN Deep Reinforcement Learning with FL
@@ 2.3 Heterogeneous Federated Learning
Existing FL approaches predominantly assume that all clients share a uniform network structure or similar data distributions. However, practical scenarios often exhibit substantial differences among clients in terms of data distributions, model structures, communication networks, and edge devices (ye2023heterogeneous). Hierarchy naturally arises due to asymmetry in demand and supply. For instance, intersections near schools may experience different peak traffic hours compared to those near sports arenas. To address FL heterogeneity, several state-of-the-art methods have been proposed. These methods, known as HFL, are categorized into data-level, model-level, and server-level approaches(ye2023heterogeneous). Among these approaches, it is crucial to identify one that can account for the heterogeneity of different intersections by dynamically grouping similar ones and training models that reflect their specific characteristics.
@@ 2.3.1 Data-level Methods
Data-level methods focus primarily on tasks related to data management, including data augmentation and anonymization techniques. Significant contributions in this area include (li2020federated) (li2019fedmd) (yoon2021fedmix). While these efforts are aimed at mitigating statistical heterogeneity between clients and enhancing data privacy, they may not be ideally suited for grouping and training models that accurately reflect the unique characteristics of diverse datasets.
@@ 2.3.2 Model-level Methods
Model-level methods focus on operations designed specifically for the model layer, including sharing partial structures and model optimization. Recent research on $\Psi$-Net (yu2022heterogeneous) and QAvg &\& PAvg (jin2022federated) optimizes policy performance in diverse environments without data sharing, although mainly for tabular methods. Per-FedAvg (fallah2020personalized) and FedPer (arivazhagan2019federated) focus on personalized federated learning, primarily for supervised learning environments. Related work in FedFormer (hebert2022fedformer) and the KT-pFL framework (zhang2021parameterized) enhances personalization through innovative model structures but with potential limitations in direct policy updates.
@@ 2.3.3 Server-level Methods
Server-level methods focus on the central server’s engagement, including tasks such as client selection, grouping, and coordination. Related works include deep reinforcement learning (DRL) on the heterogeneous training intensity assignment problem for FL (9965742) and RL client grouping (9134408), where both focus on dynamic adjustment via RL. Other works include FedCor (tang2022fedcor) and QAvg &\& PAvg (jin2022federated).
Cluster-based and optimization-based approaches further enhance client coordination and model personalization at the server-level. Cluster-based methods infer the heterogeneity of the data distribution using model information, assigning clients to clusters for model aggregation. Examples include research on FedCHAR &\& FedCHAR-DC (li2023hierarchical), the Federated Multitask Learning (FMTL) framework (sattler2020clustered), and the Iterative Federated Clustering Algorithm (IFCA) (ghosh2020efficient). Other clustering methods, such as hierarchical clustering (briggs2020federated), EM clustering (long2023multi), and KMeans (ghosh2019robust), help FL converge faster with lower communication costs. In optimization-based approaches, client aggregation is often formulated as an integer programming problem to minimize objective functions. Such methods include Ditto (li2021ditto), the APPLE framework (luo2022adapt), and FedFomo (zhang2020personalized).
@@
$t$ FL server training round
$\tau$ local training round
@@
$\mathcal{L}$ FedFomo optimization objectives
@@ 3.1 MARL
Formally, the ATSC on a road network problem is modeled as a Markov game defined by the tuple $\bigl(N,S,O,A,P,R,\gamma\bigr)$, where $N$ is the number of agents (intersections), and $S,O,A,P,R,\gamma$ denote the state space, joint observation space, action space, state transition functions, reward functions, and discount factor, respectively. Since each agent has only partially observability of the global state, the Markov game is formulated as a partially observable Markov decision process (POMDP). At each time step, each agent chooses an action based on its local policy, transitioning the environment to a new state, and generating a local reward. This cycle repeats until a terminal condition is reached, and agents refine their policies to maximize their individual rewards. However, such self-interested optimization may not directly lead to improved network-wide performance, which motivates our proposed HFRL framework in Sec. 3.2.
@@
    State ($S$). For a traffic light $n$ at round $\tau$, the local state includes lane occupancy $\bigl(o_{n}^{\tau}\bigr)$, queue length $\bigl(q_{n}^{\tau}\bigr)$, average speed $\bigl(\overline{u}_{n}^{\tau}\bigr)$, and the phase ratios $\bigl(ph_{n}^{\tau}\bigr)$ for the relevant signal phases (e.g., green, yellow, red). We follow part of the settings in Hudson’s work (hudson2022smart). Although traffic variables are aggregated across lanes, the inclusion of phase ratios preserves information about which lane groups are currently being served, allowing the agent to distinguish traffic conditions across different active phases.
    Table 3: Summary of local state features
    Feature Definition Unit Normalization Aggregation Level
    Lane Occupancy ($o_{n}^{\tau}$) Ratio of lane length occupied by vehicles across lanes controlled by traffic light $n$. $[0,1]$ (inherently normalized) Per signal (averaged over lanes)
    Queue Length ($q_{n}^{\tau}$) Ratio of lane length occupied by halted vehicles (speed $\leq 0.1$ m/s) across lanes controlled by traffic light $n$. $[0,1]$ (inherently normalized) Per signal (averaged over lanes)
    Average Speed ($\overline{u}_{n}^{\tau}$) Average vehicle speed on lanes controlled by traffic light $n$. m/s Divided by road speed limit to $[0,1]$ Per signal (aggregated across lanes)
    Phase Ratios ($ph_{n}^{\tau}$) Current-phase encoding. Ratio of signal colors (green, yellow, red) in the currently active phase of traffic light $n$. $[0,1]$ (per phase state) Per signal (current phase)
@@
    Action ($A$). To ensure consistency across agents in the federated learning setting, we define a binary action space $a_{n}^{\tau}\in\{0,1\}$. Here, $a_{n}^{\tau}=1$ indicates that traffic light $n$ will attempt to switch to the next phase at round $\tau$, while $a_{n}^{\tau}=0$ means that it remains in its current phase. The phase sequence itself follows a fixed, pre-defined transition order that ensures non-conflicting traffic flows. When the agent selects the switch action, the system advances to the next phase in this preset cycle (e.g., green → yellow → red → green), subject to minimum and maximum timing constraints between transitions (hudson2022smart).